OpenStudy (precal):

Disk Method: The bounded region formed by y=x^2, x = 3, and the x-axis is rotated about the line x = 3. Find the volume of the solid.

OpenStudy (precal):

I believe this is a dy problem

OpenStudy (precal):

limits are 0 to 9 correct?

OpenStudy (anonymous):

yes

OpenStudy (precal):

\[\pi \int\limits_{0}^{9}\left[ \sqrt{y}-3 \right]^2 dy\]

OpenStudy (precal):

did I set it up correctly?

OpenStudy (anonymous):



OpenStudy (precal):

-9pi I believe is the solution

OpenStudy (anonymous):

volumue can't be negative

OpenStudy (anonymous):

what is the radius?

OpenStudy (precal):

I dropped off the second power, but you are correct

OpenStudy (precal):

ok I got 42.4115 I just need to integrate it by hand

OpenStudy (anonymous):

your set up is correct. Technically it should be V = pi (3 - sqrt(y))^2 dy, from 0 to 9 but since the difference is squared, the answer will still come out positive

OpenStudy (precal):

Is it right - left or left - right?

OpenStudy (precal):

I have my functions switched above

OpenStudy (precal):

I believe it is right - left. I guess I still don't know my right from my left (my mistake)

OpenStudy (anonymous):

yeah, right - left

OpenStudy (precal):

no wonder I washed out of the military. haha

OpenStudy (anonymous):

:D

OpenStudy (precal):

13.5 pi or 42.411 got it Thanks